# Generative model for financial time series trained with MMD using a signature kernel

### v2

@@ 5.6 Stylised Facts
In this section, we demonstrate that the generative model is able to produce sequences that exhibit most of the behaviours in line with well known stylised facts (see e.g. [6, Section 2]) of financial time series. Recall that the training procedure involves sampling segments of the historical S&P 500 data, which is a sequence of 250 prices. We compare the statistics of the generated sequences, which have the same length as the segments, with the statistics of the segments and full sequence i.e. from 1995 to 2018. We generate sequences with the same procedure as in training, with $k=50,n=299$ in the notation of equation (22), and only use the truncated sequences, ${\mathbf{x}}_{\Theta,\text{trun}}$ in Algorithm 1, for the analysis.
As a comparison, we train a COT-GAN generative model introduced in [40] using the same data. We also use the same MA model to generate the noise input to the model, which has the same architecture including the number of noise dimensions. To ensure a fair comparison, we also use the same noise sequence for both models to generate the samples of the same length. We modify the code from the authors’ repository available at https://github.com/tianlinxu312/cot-gan for our specific use case.
@@ 5.6.1 Moments of Returns
In addition, we also include the statistics of the moments for both our MMD based generator and the COT-GAN generator in Table 6. The table shows the mean and standard deviation of the distribution of the moments across 10,000 samples. The COT-GAN samples have a wider distribution of the first moment compared to the MMD generator. This is also evident when we look at the distribution of the end points of the generated segments.
@@
Figure 10 shows the density plot of the generated sequences’ end points with the initial points normalised to 1. Note that we truncated the x-axis for better visualisation as 35 of the COT-GAN samples had end points scattered between 2.5 and 5.6. The end points of the MMD generated sequences are more aligned with the S&P 500 segments than the COT-GAN generated sequences. The presence of more extreme returns explains the higher 2nd moment and lower 4th moment of the COT-GAN generated sequences due to a fatter middle. The extreme values of the COT-GAN generated sequences are overly pronounced and are not a realistic reflection of the training data.
Figure 10: Density plot of the end points of the generated sequences. Initial values are normalised to start from 1.0 for all sequences. The x-axis is truncated (due to more extreme vales from the COT-GAN generated sequences) for better visualisation.
@@ 5.6.2 Autocorrelation in Returns
In Figure 11(a), the statistics are consistent across the 3 types of sequences where it shows a lack of linear autocorrelation. However, returns are certainly not independent due to the phenomenon of volatility clustering. There are different manifestations of volatility clustering and one of them is the autocorrelation of the squared or absolute returns. Figures 11(b) and 11(c) show the autocorrelation of the squared and absolute returns respectively. While both the S&P 500 segments and full sequence exhibit positive and slow decaying autocorrelation with respect to an increasing time lag, the full sequence shows a more pronounced effect. The generated sequences are a closer match to the sampled sequences than to the full sequence as expected due to the training procedure. The COT-GAN generated sequences show a similar pattern to the MMD generated sequences but are omitted for brevity [^14] See footnote 12.
@@
(c) Autocorrelation of absolute returns
Figure 11: Linear and non-linear autocorrelations of daily log returns of varying lags. Left, middle and right figure show autocorrelation coefficient of the returns, squared returns and absolute returns, respectively. The blue and orange lines are the mean of the autocorrelation coefficient across the generated and sampled sequences, respectively while the shaded region shows the range of the mean absolute deviation. The red line is the autocorrelation coefficient of the full S&P 500 sequence.
@@ 5.6.3 Gain/loss Asymmetry and Leverage Effect
The negative skew in the observed returns can also be viewed through different lenses. The gain/loss asymmetry is the phenomenon that there are more large negative returns observed than large positive returns (see e.g.[6]). Figure 12(a) shows the expected ratio of positive returns (y-axis) given that the absolute value of the return exceeds a certain threshold (x-axis). For the threshold, we only use values up to 0.0237 which corresponds to the 95th percentile of the returns found in S&P 500 since there are limited samples beyond this point. On this measure, the MMD generated sequences are more aligned to the S&P 500 segments displaying a similar level of asymmetry. However, the COT-GAN generated sequences have a return distribution that is more symmetric.
The leverage effect is the observation that the volatility of the returns is negatively correlated with the returns (see e.g.[6]). We can quantify this by calculating the correlation between the returns and the squared returns in the subsequent days as shown in Figure 12(b). On this measure, we see that the generated sequences only show a weak negative correlation in the immediate subsequent day before it quickly decays towards zero. This is in contrast to the S&P 500 segments and full sequence which shows a more pronounced negative correlation that is slower to decay. We omit the COT-GAN generated sequences for brevity as they show a similar pattern to the MMD generated sequences [^15] See footnote 12.
@@
Figure 12: The left plot shows the gain/loss asymmetry of the daily log returns. The y-axis is the expected ratio of positive returns given that the return exceeds a certain threshold on the x-axis. The right plot shows the correlation between the daily log returns and the squared returns of the subsequent day.
@@ 6 Conclusion and Future Work
This paper introduces an approach for generating realistic synthetic financial time series data using MMD with a signature kernel. The use of a moving average (MA) model to vary the variance of the noise input to the generative model aids the performance. In particular, the model demonstrates the ability to capture key stylised facts observed in financial markets.

### v3

@@ 5.6 Stylised Facts
In this section, we demonstrate that the generative model is able to produce sequences that exhibit most of the behaviours in line with well known stylised facts (see e.g. [12, Section 2]) of financial time series. Recall that the training procedure involves sampling segments of the historical S&P 500 data, which is a sequence of 250 prices. We compare the statistics of the generated sequences, which have the same length as the segments, with the statistics of the segments and full sequence i.e. from 1995 to 2018. We generate sequences with the same procedure as in training, with $k=50,n=299$ in the notation of equation (22), and only use the truncated sequences, ${\mathbf{x}}_{\Theta,\text{trun}}$ in Algorithm 1, for the analysis.
As a comparison, we train a COT-GAN and Conditional Sig-WGAN (C-Sig-WGAN) generative models introduced in [59] and [37] respectively using the same data. We also use the same MA model to generate the noise input to the COT-GAN model, which has the same architecture including the number of noise dimensions. To ensure a fair comparison, we also use the same noise sequence for both models to generate the samples of the same length. The Conditional Sig-WGAN model is set up with the original optimised specification in [37] except with the signature truncation order increased to 5 which was the limit of the GPU memory available for the algorithm to run. We modify the code from the authors’ repository available at https://github.com/tianlinxu312/cot-gan for COT-GAN and https://github.com/SigCGANs/Conditional-Sig-Wasserstein-GANs for Conditional Sig-WGAN for our specific use case.
@@ 5.6.1 Moments of Returns
In addition, we also include the statistics of the moments for both our MMD based generator, the COT-GAN generator and the C-Sig-WGAN generator in Table 6. The table shows the mean and standard deviation of the distribution of the moments across 10,000 samples. In general, the models more or less capture the characteristics of the S&P 500. One deviation is in the samples from the C-Sig-WGAN which on average produces a slightly positive skew. The COT-GAN samples have a much wider distribution of the first moment compared to the other models and the S&P 500 segments. This is also evident when we look at the distribution of the end points of the generated segments.
@@
Figure 10 shows the density plot of the generated sequences’ end points with the initial points normalised to 1. Note that we truncated the x-axis for better visualisation as 35 of the COT-GAN samples had end points scattered between 2.5 and 5.6. The end points of the MMD generated sequences are more aligned with the S&P 500 segments than the COT-GAN generated sequences. The presence of more extreme returns explains the higher 2nd moment and lower 4th moment of the COT-GAN generated sequences due to a fatter middle. The extreme values of the COT-GAN generated sequences are overly pronounced and are not a realistic reflection of the training data. For the C-Sig-WGAN generated sequences, it is closer to the training data but the MMD generated sequences have a better overall fit.
(a) MMD
(b) COT-GAN
(c) C-Sig-WGAN
Figure 10: Density plot of the end points of the generated sequences. Initial values are normalised to start from 1.0 for all sequences. The x-axis is truncated for COT-GAN (due to more extreme vales). The S&P 500 histogram is constructed based on the training data construction method which are the same for MMD and COT-GAN but slightly different for C-Sig-WGAN which has shorter lengths.
@@ 5.6.2 Autocorrelation in Returns
In Figure 11(a), the statistics are consistent across the 3 types of sequences where it shows a lack of linear autocorrelation. However, returns are certainly not independent due to the phenomenon of volatility clustering. There are different manifestations of volatility clustering and one of them is the autocorrelation of the squared returns. Figure 11(b) shows the autocorrelation of the squared returns. While both the S&P 500 segments and full sequence exhibit positive and slow decaying autocorrelation with respect to an increasing time lag, the full sequence shows a more pronounced effect. The generated sequences are a closer match to the sampled sequences than to the full sequence as expected due to the training procedure.
@@
Figure 11: Linear and non-linear autocorrelations of daily log returns of varying lags. Left and right figures show autocorrelation coefficient of the returns and squared returns, respectively. The blue and orange lines are the mean of the coefficient across the generated and sampled sequences, respectivel, while the shaded region shows the range of the mean absolute deviation. The red line is the coefficient of the full S&P 500 sequence.
Comparing the MMD generator with the COT-GAN and C-Sig-WGAN generators in Figure 12, we see that the MMD and COT-GAN generators produce similar autocorrelation of the squared returns in Figure 12(a). The C-Sig-WGAN generated sequences exhibit a much faster and smoother decay in the autocorrelation of the squared returns as shown in Figure 12(b). We omit the comparison for linear autocorrelation of the returns as the COT-GAN and C-Sig-WGAN generators produce similar results. [^14] See footnote 12.
(a) MMD vs COT-GAN for squared returns
(b) MMD vs C-Sig-WGAN for squared returns
Figure 12: Autocorrelation of squared returns for MMD, COT-GAN and C-Sig-WGAN. Lines and shaded regions are the same as in Figure 11.
@@ 5.6.3 Gain/loss Asymmetry and Leverage Effect
The negative skew in the observed returns can also be viewed through different lenses. The gain/loss asymmetry is the phenomenon that there are more large negative returns observed than large positive returns (see e.g.[12]). Figure 13(a) shows the expected ratio of positive returns (y-axis) given that the absolute value of the return exceeds a certain threshold (x-axis). For the threshold, we only use values up to 0.0237 which corresponds to the 95th percentile of the returns found in S&P 500 since there are limited samples beyond this point. On this measure, the MMD generated sequences are more aligned to the S&P 500 segments displaying a similar level of asymmetry. However, the COT-GAN generated sequences have a return distribution that is more symmetric. For the C-Sig-WGAN generated sequences, the asymmetry more or less follows the same pattern as the S&P 500 but is much more pronounced until it gets closer to the 95th percentile.
The leverage effect is the observation that the volatility of the returns is negatively correlated with the returns (see e.g.[12]). We can quantify this by calculating the correlation between the returns and the squared returns in the subsequent days as shown in Figure 13(b). On this measure, we see that the MMD generated sequences only show a weak negative correlation in the immediate subsequent day before it quickly decays towards zero. This is in contrast to the S&P 500 segments and full sequence which shows a more pronounced negative correlation that is slower to decay. The COT-GAN generated sequences show a similar pattern to the MMD generated sequences while the C-Sig-WGAN exhibits a pronounced negative correlation in the immediate subsequent day before it quickly decays towards zero.
@@
(c) MMD vs COT-GAN for leverage effect
(d) MMD vs C-Sig-WGAN for leverage effect
Figure 13: The top left plot shows the gain/loss asymmetry of the daily log returns. The y-axis is the expected ratio of positive returns given that the return exceeds a certain threshold on the x-axis. The other three plots show the correlation between the daily log returns and the squared returns of the subsequent day with lines and shaded regions being the same as in Figure 11.
@@ 6 Conclusion and Future Work
This paper introduces an approach for generating realistic synthetic financial time series data using MMD with a signature kernel. The use of a moving average (MA) model to vary the variance of the noise input to the generative model aids the performance. In particular, the model demonstrates the ability to capture key stylized facts observed in financial markets.